OpenStudy (anonymous):

Find the difference. (16x3 + 2x2) - 4x3 = Select one of the options below as your answer: A. 14x3 B. 12x3 - 2x2 C. 12x3 + 2x2 D. 16x3 - 2x2

OpenStudy (munish):

(16x3+2x2)-4x3 =16x3+2x2-4x3 =16x3-4x3+2x2 =12x3+2x2 so c)